OpenStudy (anonymous):

evaluate a^-2b^0 for a =3 and b=4

OpenStudy (gw2011):

If you mean: (a^-2)(b^0), then you get: (3^-2)(4^0)=[(1/3)^2][1]=1/9

OpenStudy (anonymous):

\[3^{-2} 4^0\] \[3^{-2}\times 1\]

OpenStudy (anonymous):

\[3^{-2}=\frac{1}{3}^2=\frac{1}{9}\]